# ROVER: A Multi-Season Dataset for Visual SLAM

### v1

@@ Abstract
Robust Simultaneous Localization and Mapping (SLAM) is a crucial enabler for autonomous navigation in natural, unstructured environments such as parks and gardens. However, these environments present unique challenges for SLAM due to frequent seasonal changes, varying light conditions, and dense vegetation. These factors often degrade the performance of visual SLAM algorithms originally developed for structured urban environments. To address this gap, we present ROVER, a comprehensive benchmark dataset tailored for evaluating visual SLAM algorithms under diverse environmental conditions and spatial configurations. We captured the dataset with a robotic platform equipped with monocular, stereo, and RGB-D cameras, as well as inertial sensors. It covers 39 recordings across five outdoor locations, collected through all seasons and various lighting scenarios, i.e., day, dusk, and night with and without external lighting. With this novel dataset, we evaluate several traditional and deep learning-based SLAM methods and study their performance in diverse challenging conditions. The results demonstrate that while stereo-inertial and RGB-D configurations generally perform better under favorable lighting and moderate vegetation, most SLAM systems perform poorly in low-light and high-vegetation scenarios, particularly during summer and autumn. Our analysis highlights the need for improved adaptability in visual SLAM algorithms for outdoor applications, as current systems struggle with dynamic environmental factors affecting scale, feature extraction, and trajectory consistency. This dataset provides a solid foundation for advancing visual SLAM research in real-world, natural environments, fostering the development of more resilient SLAM systems for long-term outdoor localization and mapping. The dataset and the code of the benchmark are available under https://iis-esslingen.github.io/rover.
@@ I Introduction
Simultaneous Localization and Mapping (SLAM) is a foundational technology in mobile robotics, enabling autonomous systems to construct a map of an unknown environment while simultaneously determining their location within it. This capability is crucial for robots to navigate effectively in dynamic, unstructured environments where external positioning systems, such as Global Navigation Satellite Systems (GNSS), may be unreliable or unavailable [1, 2, 3, 4]. In scenarios such as precision agriculture, forestry, and rugged outdoor environments, high-precision localization is essential for automation, particularly under challenging and constantly changing environmental conditions [5].
@@
Visual SLAM, in particular, has emerged as a key approach due to its ability to leverage visual data from cameras, making it a cost-effective and versatile solution [6]. However, most existing visual SLAM datasets are centered around structured environments such as indoor spaces [7] or urban areas [8], which lack the unique challenges presented by natural outdoor environments. Natural settings introduce a wide range of unpredictable factors, seasonal variations, fluctuating daylight, dynamic weather, and uneven terrain that affect long-term, robust localization and mapping. Visual SLAM systems must be resilient to both static and dynamic changes in these environments to achieve reliable performance over time [9, 10, 11]. Current datasets often lack comprehensive coverage of environmental conditions in semi-structured and natural outdoor settings, such as parks and gardens, where factors like lighting, seasonal changes, and weather are less commonly represented [12]. This creates a notable gap in visual SLAM research, particularly for systems using diverse camera setups including monocular, stereo, and RGB-D configurations. Effective assessment of environmental impacts on localization and mapping, especially for multi-session mapping across varying conditions, remains underexplored and is crucial for advancing visual SLAM systems in such applications [13]. While seasonal changes are occasionally captured in agricultural datasets [14, 15, 16, 17, 18, 19], they are rarely present in datasets of natural environments such as parks and gardens, where visual complexity varies widely.
@@
In addition to the dataset, this work provides a benchmark of traditional and deep learning-based visual SLAM methods, evaluating their performance across these challenging outdoor environments. These methods are applied across various sensor modalities, utilizing monocular, stereo, and RGB-D approaches in combination with inertial data, allowing for a comprehensive comparison between traditional visual SLAM algorithms and data-driven deep learning approaches. The goal is to provide a detailed evaluation of how well these methods perform in real-world, natural outdoor environments that pose significant challenges for localization and mapping.
@@
  • A millimeter-accurate localization benchmark dataset capturing extensive multi-session trajectories across diverse natural outdoor environments, utilizing a multi-sensor platform with various camera types and IMUs for robust evaluation of different SLAM configurations.
@@ II Related Work
To support the development of SLAM algorithms, diverse datasets are essential to capture the variety of conditions these systems must handle across different applications. SLAM methods often vary in their approach, depending on the target application, environment, and sensor setup, each requiring a unique range of data for thorough evaluation. In this study, we focus on datasets that capture natural settings using wheeled or legged robots, cars, or handheld devices, as these environments present unique challenges such as seasonal variation, dynamic weather, and complex lighting scenarios. In the following sections, we discuss related outdoor datasets that provide diverse evaluation conditions for SLAM. These include datasets focused on both structured and unstructured settings such as urban and forested areas, as well as agricultural and garden environments. Table I provides an overview of these related datasets, detailing each dataset’s platform, environment type, diversity of scenarios, sensor configuration, and availability of ground truth data. This table also highlights how our dataset differs by focusing specifically on various natural outdoor settings that are underrepresented in existing datasets.
@@
Trajectory Weather Lighting Season RGB Stereo RGB-D IMU
@@
FoMo [35] UGV Forest x x x x x x RTK-GNSS
@@
Rally Estonia [40] Car Forest, Rural x x x x x GNSS
@@
UTIAS Long-Term [43] UGV Off-Road x x x x x x VT&R
@@
  • Abbreviations: RTK - Real-Time Kinematic, INS - inertial Navigation System, VT&R - Visual Teach and Repeat [54], ICP - Iterative Closest Point [55]
@@
The DiTer [47] and DiTer++ [48] datasets focus on park-like and forested areas, using legged robots to capture various terrains, including sandy roads and slopes. DiTer++ expands on the original dataset with multi-robot setups, heterogeneous LiDARs, and day-night sequences, making it more versatile for SLAM evaluation. The Newer College [49] and Newer College Multi-Camera [50] datasets are collected on the campus of New College, Oxford. The original dataset, captured with a handheld system, includes structured buildings, vegetation, and open spaces, while the extended version introduces a multi-camera system with fisheye lenses, addressing challenges like abrupt lighting changes and textureless surfaces. The BotanicGarden dataset [51], captured in a diverse botanical garden using a UGV platform, features thick woods, riversides, narrow trails, and open grasslands, providing varied scenarios for testing SLAM in natural and semi-structured environments. The EDEN dataset [52] is a synthetic multimodal dataset with over 100 garden models, offering rich variations in landscapes and vegetation. It provides data under different weather and lighting conditions, inspired by the TB-Places dataset [53], which is designed for visual place recognition in garden environments. TB-Places, captured with a lawn mower, include diverse terrain types and obstacles, filling a critical gap in natural environment place recognition data.
Existing datasets often lack comprehensive coverage of diverse environmental conditions, particularly in garden and park-like settings, where variations in lighting, seasons, and weather are underrepresented, refer to Table I. Furthermore, there is a significant need for datasets that specifically assess the impact of these environmental factors on localization and mapping, particularly for visual SLAM systems using various camera types, such as monocular, stereo, and RGB-D, with the potential for use in multi-session mapping studies. To address these gaps, our dataset offers a high-precision, long-term benchmark, capturing consistent routes across five distinct natural outdoor environments under a wide range of conditions, including varied weather and lighting scenarios across all seasons. The dataset also addresses challenging real-world factors, such as dynamic illumination changes, reflective and textureless surfaces, motion blur, occlusions, and uneven terrain, providing a comprehensive resource for evaluating and advancing visual SLAM algorithms.
III Data Recording System
III-A Robotic Platform
@@ III-B1 Cameras
Cameras are the central component of the dataset, providing diverse perspectives and depth of information. The cameras are mounted at the front of the sensor platform and include monocular, stereo, and RGB-D cameras, each selected to capture different environmental aspects, see Figure 2.
@@
The choice of cameras reflects a balance between cost and functionality. Monocular cameras such as the PiCam provide standard RGB images, while stereo and RGB-D cameras such as the D435i and T265 enhance depth perception and spatial awareness. These cameras also feature different fields of view, from wide-angle to fisheye lenses, to maximize environmental coverage. Figure 3 highlights the diverse perspectives captured by each camera.
@@ III-C2 Time Synchronization
The time synchronization across the entire system is crucial for ensuring the temporal alignment of data collected from the sensors and the ground truth system. This synchronization is achieved through a GNSS signal provided by a u-blox GPS receiver, which serves as the primary time reference for the system. Both Pi devices are connected to a u-blox C099-F9P Application Board, featuring a u-blox ZED-F9P-01B module that receives accurate GNSS signals. On the sensor platform, one of the Pis serves as the NTP (Network Time Protocol) server and is connected via Ethernet to a network switch. This switch links the NTP server to the other computing devices, ensuring that all sensor computing units share a common time base. By leveraging the GNSS signal and NTP synchronization, the system achieves nanosecond-level accuracy, allowing the NTP server to synchronize the NTP clients on these devices and provide precise timestamps for data collection.
@@ IV ROVER Dataset
Designed to present a diverse set of environmental conditions, this dataset provides a comprehensive basis for evaluating the robustness of SLAM algorithms. With varied locations, seasonal shifts, and environmental factors such as weather and lighting, it enables a thorough assessment of SLAM adaptability in natural and semi-structured settings. The following subsections detail the recording locations, their defining characteristics, and the methodology behind data collection, highlighting the conditions under which the data was acquired.
@@ IV-A Locations
(a) Park
(b) Campus Small
(c) Campus Large
(d) Garden Small
(e) Garden Large
@@ V Visual SLAM Benchmark
TABLE V: Overview of SLAM methods, supported sensor configurations, and algorithmic components.
SLAM Mono Mono- Stereo Stereo- RGB-D RGB-D- Frontend Estimation Backend Loop
@@
V-A Visual SLAM Methods
We explore a variety of visual and visual-inertial SLAM methods, employing both traditional and deep learning-based approaches, to address different aspects of visual localization and mapping. The methods differ in their design, ranging from feature-based and semi-direct techniques to deep learning methods that utilize dense direct approaches with optical flow estimation. Additionally, some methods incorporate loop closing for long-term consistency, while others do not. The approaches also vary in estimation techniques, utilizing either filter-based or optimization-based backends to support different sensor modalities, such as monocular, stereo, and RGB-D, with or without inertial data integration. Table V provides a detailed overview of the capabilities and configurations of the SLAM methods considered in this study.
@@ V-B Evaluation Methodology
Once the valid trajectories are identified, alignment is performed to accurately compare the estimated and ground truth trajectories. Since the trajectories may exist in different coordinate systems, we employ Umeyama’s method [70] to find the optimal transformation that maps the estimated trajectory onto the ground truth, yielding a least-squares fit. Following Zhang and Scaramuzza [71], SLAM methods rely solely on monocular systems, the trajectories are scaled relative to the ground truth using a SIM(3) transformation to account for scale ambiguity. In contrast, for other modalities such as stereo, RGB-D, or those incorporating inertial data, where the scale is observable, an SE(3) transformation (including only rotation and translation) is applied. This alignment step ensures that the subsequent evaluation measures the true deviation from the ground truth.
Finally, we calculate the root mean squared error (RMSE) of the Absolute Trajectory Error (ATE) as the primary metric to quantify the global accuracy of the SLAM system. The ATE captures the overall difference between the aligned estimated trajectory and the ground truth, following the methodology outlined in [72]. To ensure temporal consistency, linear interpolation is applied to synchronize the timestamps between the estimated and ground truth data points before computing the RMSE. The average RMSE ATE is calculated using the open-source Evo toolbox [73]. All experiments were conducted on an AMD Ryzen 9 7950X3D with 64 GB of RAM and an NVIDIA RTX A6000 within a Docker container based on Ubuntu 20.04 and ROS Noetic. We utilized open-source implementations of each SLAM algorithm and applied pretrained weights for DROID-SLAM, DPVO, and DPV-SLAM to ensure reproducibility and consistency in the evaluation.

### v2

@@ Abstract
Robust Simultaneous Localization and Mapping (SLAM) is a crucial enabler for autonomous navigation in natural, semi-structured environments such as parks and gardens. However, these environments present unique challenges for SLAM due to frequent seasonal changes, varying light conditions, and dense vegetation. These factors often degrade the performance of visual SLAM algorithms originally developed for structured urban environments. To address this gap, we present ROVER, a comprehensive benchmark dataset tailored for evaluating visual SLAM algorithms under diverse environmental conditions and spatial configurations. We captured the dataset with a robotic platform equipped with monocular, stereo, and RGBD cameras, as well as inertial sensors. It covers 39 recordings across five outdoor locations, collected through all seasons and various lighting scenarios, i.e., day, dusk, and night with and without external lighting. With this novel dataset, we evaluate several traditional and deep learning-based SLAM methods and study their performance in diverse challenging conditions. The results demonstrate that while stereo-inertial and RGBD configurations generally perform better under favorable lighting and moderate vegetation, most SLAM systems perform poorly in low-light and high-vegetation scenarios, particularly during summer and autumn. Our analysis highlights the need for improved adaptability in visual SLAM algorithms for outdoor applications, as current systems struggle with dynamic environmental factors affecting scale, feature extraction, and trajectory consistency. This dataset provides a solid foundation for advancing visual SLAM research in real-world, semi-structured environments, fostering the development of more resilient SLAM systems for long-term outdoor localization and mapping. The dataset and the code of the benchmark are available under https://iis-esslingen.github.io/rover.
@@ I Introduction
Simultaneous Localization and Mapping (SLAM) is a foundational technology in mobile robotics, enabling autonomous systems to construct a map of an unknown environment while simultaneously determining their location within it. This capability is crucial for robots to navigate effectively in dynamic, unstructured environments where external positioning systems, such as Global Navigation Satellite Systems (GNSS), may be unreliable or unavailable. In scenarios such as precision agriculture, forestry, and rugged outdoor environments, high-precision localization is essential for automation, particularly under challenging and constantly changing environmental conditions [1].
@@
Visual SLAM, in particular, has emerged as a key approach due to its ability to leverage visual data from cameras, making it a cost-effective and versatile solution [2]. However, most existing visual SLAM datasets are centered around structured environments such as indoor spaces [3] or urban areas [4], which lack the unique challenges presented by natural outdoor environments. Natural settings introduce a wide range of unpredictable factors, seasonal variations, fluctuating daylight, dynamic weather, and uneven terrain that affect long-term, robust localization and mapping. Visual SLAM systems must be resilient to both static and dynamic changes in these environments to achieve reliable performance over time [5]. While certain existing datasets capture multi-season data in agricultural contexts (e.g., [6, 7, 8, 9, 10, 11]) or forest and off-road environments (e.g., [12, 13]), few focus on semi-structured, garden-like outdoor settings. These environments, such as parks or suburban lawns, introduce a distinct mix of man-made structures and natural features, along with moderate terrain variations, dense vegetation, and varying lighting conditions. Effective assessment of environmental impacts on localization and mapping, especially for multi-session mapping across varying conditions, remains underexplored and is crucial for advancing visual SLAM systems [14]. Consequently, there remains a notable gap in visual SLAM research for diverse camera setups (monocular, stereo, RGBD) under long-term seasonal and lighting changes in such semi-structured scenarios.
@@
In addition to the dataset, this work provides a benchmark of traditional and deep learning-based visual SLAM methods, evaluating their performance across these challenging outdoor environments. These methods are applied across various sensor modalities, utilizing monocular, stereo, and RGBD approaches in combination with inertial data, allowing for a comprehensive comparison between traditional visual SLAM algorithms and data-driven deep learning approaches. The goal is to provide a detailed evaluation of how well these methods perform in semi-structured outdoor environments that pose significant challenges for localization and mapping.
@@
  • A millimeter-accurate localization benchmark dataset capturing extensive multi-session trajectories across diverse outdoor environments, utilizing a multi-sensor platform with various camera types and IMUs for robust evaluation of different SLAM configurations.
@@ II Related Work
To support the development of SLAM algorithms, diverse datasets are essential to capture the variety of conditions these systems must handle across different applications. SLAM methods often vary in their approach, depending on the target application, environment, and sensor setup, each requiring a unique range of data for thorough evaluation. In this study, we focus on datasets that capture natural settings using wheeled or legged robots, cars, or handheld devices, as these environments present unique challenges such as seasonal variation, dynamic weather, and complex lighting scenarios. Note that numerous large-scale autonomous driving datasets also address seasonal changes in urban environments [17, 18, 19, 20, 21], but we do not discuss them in detail here as our emphasis is on unstructured and semi-structured, natural outdoor settings that are less represented in these urban-focused datasets. In the following sections, we discuss related outdoor datasets that provide diverse evaluation conditions for SLAM. These include datasets focused on both structured and unstructured settings such as urban and forested areas, as well as agricultural and garden environments. Table I provides an overview of these related datasets, detailing each dataset’s platform, environment type, diversity of scenarios, sensor configuration, and availability of ground truth data. This table also highlights how our dataset differs by focusing specifically on various natural outdoor settings that are underrepresented in existing datasets.
@@
Trajectory Weather Lighting Season Mono Stereo RGBD IMU
@@
FoMo [12] UGV Forest x x x x x x PPK-GNSS
@@
Rally Estonia [39] Car Forest, Rural (x) x x x x GNSS
@@
UTIAS Long-Term [13] UGV Campus, Off-Road x x x x x x VT&R
@@
  • Abbreviations: RTK - Real-Time Kinematic, PPK - Post-Processing Kinematic, INS - Inertial Navigation System, VT&R - Visual Teach and Repeat [52], ICP - Iterative Closest Point [53]
@@
The DiTer [45] and DiTer++ [46] datasets focus on park-like and forested areas, using legged robots to capture various terrains, including sandy roads and slopes. DiTer++ expands on the original dataset with multi-robot setups, heterogeneous LiDARs, and day-night sequences, making it more versatile for SLAM evaluation. The Newer College [47] and Newer College Multi-Camera [48] datasets are collected on the campus of New College, Oxford. The original dataset, captured with a handheld system, includes structured buildings, vegetation, and open spaces, while the extended version introduces a multi-camera system with fisheye lenses, addressing challenges like abrupt lighting changes and textureless surfaces. The BotanicGarden dataset [49], captured in a diverse botanical garden using a UGV platform, features thick woods, riversides, narrow trails, and open grasslands, providing varied scenarios for testing SLAM in natural and semi-structured environments. The EDEN dataset [50] is a synthetic multimodal dataset with over 100 garden models, offering rich variations in landscapes and vegetation. It provides data under different weather and lighting conditions, inspired by the TB-Places dataset [51], which is designed for visual place recognition in garden environments. TB-Places, captured with a lawn mower, includes diverse terrain types and obstacles, filling a critical gap in natural environment place recognition data.
@@ III-B1 Cameras
Cameras are the central component of the dataset, providing diverse perspectives and depth of information. The cameras are mounted at the front of the sensor platform and include monocular, stereo, and RGBD cameras, each selected to capture different environmental aspects, see Figure 2.
@@
The choice of cameras reflects a balance between cost and functionality. Monocular cameras such as the PiCam provide standard RGB images, while stereo and RGBD cameras such as the D435i and T265 enhance depth perception and spatial awareness. These cameras also feature different fields of view, from wide-angle to fisheye lenses, to maximize environmental coverage. Figure 3 highlights the diverse perspectives captured by each camera.
@@ III-C2 Time Synchronization
The time synchronization across the entire system is crucial for ensuring the temporal alignment of data collected from the sensors and the ground truth system. This synchronization is achieved through a GNSS signal provided by a u-blox GPS receiver, which serves as the primary time reference for the system. Both Pi devices are connected to a u-blox C099-F9P Application Board, featuring a u-blox ZED-F9P-01B module that receives accurate GNSS signals. On the sensor platform, one of the Pis serves as the NTP (Network Time Protocol) server and is connected via Ethernet to a network switch. This switch links the NTP server to the other computing devices, ensuring that all sensor computing units share a common time base. By leveraging the GNSS signal and NTP synchronization, the system achieves millisecond-level accuracy, allowing the NTP server to synchronize the NTP clients on these devices and provide precise timestamps for data collection.
@@ IV ROVER Dataset
Designed to present a diverse set of environmental conditions, this dataset provides a comprehensive basis for evaluating the robustness of SLAM algorithms. With varied locations, seasonal shifts, and environmental factors such as weather and lighting, it enables a thorough assessment of SLAM adaptability in natural and semi-structured settings. The following subsections detail the recording locations, their defining characteristics, and the methodology behind data collection, highlighting the conditions under which the data was acquired, and also discuss the organization of the dataset.
@@ IV-A Locations
(a) Park
(b) Campus Small
(c) Campus Large
(d) Garden Small
(e) Garden Large
@@ IV-B Recording Overview
IV-C Dataset Organisation
The dataset includes recordings from multiple locations and dates (see Table IV). For each combination of <location> and <date>, the data is organized into folders following the hierarchy illustrated in Figure 7. Inside each <date> folder, we provide separate directories for the D435i, the T265, and the PiCam, as well as a folder for VN100 IMU data and a file for the 3 DOF ground truth trajectory (groundtruth.txt).
All RGB and depth images for the D435i are named by their Unix timestamp (<timestamp>) and saved as a PNG file. We include rgb.txt and depth.txt files following the format proposed by [55], where each line contains a timestamp and an image file path to facilitate easy parsing by common SLAM frameworks. IMU data from the D435i is stored in imu.txt, containing the 6 DoF measurements (timestamped linear acceleration and angular velocity). The T265 is organized analogous to the D435i. Two image directories (cam_left and cam_right) hold stereo PNG files, with matching cam_left.txt and cam_right.txt for timestamp-to-file mapping. The T265’s IMU measurements are stored in imu.txt, with the same format as the D435i. Similarly, the PiCam data is in the picam directory, with RGB images named by their <timestamp>, plus a file rgb.txt for time-indexed image paths. The vn100 folder contains a single imu.txt file that logs the timestamped 9 DoF (accelerometer, gyroscope, magnetometer) readings of the VN100 sensor. All calibration files are stored in the calibration directory, with one YAML file per sensor (e.g., realsense_D435i.yaml). These files contain the intrinsic parameters for each sensor; for the D435i and T265, they also include the extrinsic calibration to the internal IMU. In addition, all cameras are calibrated with respect to both the VN100 and the prism.
In addition to the raw file-based data, we provide a Python script to convert each folder’s contents into rosbags. Table V outlines the ROS topic structure, message types, and approximate publish rates for each sensor. This enables users to integrate smoothly with ROS-based SLAM and perception frameworks.
[…]
.1 /ROVER. .2 ¡location¿. .3 ¡date¿. .4 realsense_D435i. .5 rgb. .6 ¡timestamp¿.png. .5 depth. .6 ¡timestamp¿.png. .5 rgb.txt. .5 depth.txt. .5 imu. .6 imu.txt. .4 realsense_T265. .5 cam_left. .6 ¡timestamp¿.png. .5 cam_right. .6 ¡timestamp¿.png. .5 cam_left.txt. .5 cam_right.txt. .5 imu. .6 imu.txt. .4 picam. .5 rgb. .6 ¡timestamp¿.png. .5 rgb.txt. .3 vn100. .4 imu.txt. .3 groundtruth.txt. .2 calibration. .3 realsense_D435i.yaml. .3 realsense_T265.yaml. .3 picam.yaml. .3 vn100.yaml.
Figure 7: Folder structure of the ROVER dataset.
TABLE V: Rosbag overview: topics, message types, and publish rates.
Sensor ROS Topic ROS Message Type Publish Rate (hz)
Intel RealSense D435i /d435i/rgb_image sensor_msgs/Image 30
/d435i/depth_image sensor_msgs/Image 30
/d435i/imu sensor_msgs/Imu 300
Intel RealSense T265 /t265/image_left sensor_msgs/Image 30
/t265/image_left sensor_msgs/Image 30
/t265/imu sensor_msgs/Imu 265
PiCam /pi_cam/rgb_image sensor_msgs/Image 30
VN100 /vn100/imu sensor_msgs/Imu 65
@@ V Visual SLAM Benchmark
V-A Visual SLAM Methods
We explore a variety of visual and visual-inertial SLAM methods, employing both traditional and deep learning-based approaches, to address different aspects of visual localization and mapping. The methods differ in their design, ranging from feature-based and semi-direct techniques to deep learning methods that utilize dense direct approaches with optical flow estimation. Additionally, some methods incorporate loop closing for long-term consistency, while others do not. The approaches also vary in estimation techniques, utilizing either filter-based or optimization-based backends to support different sensor modalities, such as monocular, stereo, and RGBD, with or without inertial data integration. Table VI provides a detailed overview of the capabilities and configurations of the SLAM methods considered in this study.
TABLE VI: Overview of SLAM methods, supported sensor configurations, and algorithmic components.
SLAM Mono Mono- Stereo Stereo- RGBD RGBD- Frontend Estimation Backend Loop
@@ V-B Evaluation Methodology
Once the valid trajectories are identified, alignment is performed to accurately compare the estimated and ground truth trajectories. Since the trajectories may exist in different coordinate systems, we employ Umeyama’s method [69] to find the optimal transformation that maps the estimated trajectory onto the ground truth, yielding a least-squares fit. Following Zhang and Scaramuzza [70], SLAM methods rely solely on monocular systems, the trajectories are scaled relative to the ground truth using a SIM(3) transformation to account for scale ambiguity. In contrast, for other modalities such as stereo, RGBD, or those incorporating inertial data, where the scale is observable, an SE(3) transformation (including only rotation and translation) is applied. This alignment step ensures that the subsequent evaluation measures the true deviation from the ground truth.
Finally, we calculate the root mean squared error (RMSE) of the Absolute Trajectory Error (ATE) and the RMSE of the Relative Pose Error (RPE) per meter as the primary metrics to quantify the global and local accuracy of the SLAM system, following the methodology outlined in [55]. To ensure temporal consistency, linear interpolation is applied to synchronize the timestamps between the estimated and ground truth data points before computing the RMSE. The RMSE ATE and RPE are calculated using the open-source Evo toolbox [71]. In addition, we calculate the success rate (SR), which serves as a key indicator of a method’s robustness by reflecting the ratio of valid trajectories to the total number of trials for each SLAM configuration. All experiments were conducted on an AMD Ryzen 9 7950X3D with 64 GB of RAM and an NVIDIA RTX A6000 within a Docker container based on Ubuntu 20.04 and ROS Noetic. We utilized open-source implementations of each SLAM algorithm and applied pretrained weights for DROID-SLAM, DPVO, and DPV-SLAM to ensure reproducibility and consistency in the evaluation.